# Life-Code: Central Dogma Modeling with Multi-Omics Sequence Unification

### v2

@@ 1 Introduction
Figure 1: Overview of the Central Dogma Data Flow with biological language models (LM) severally proposed in each omics.
The advent of large-scale biological datasets has sparked extensive interest in sequence modeling across multiple omics domains—ranging from DNA and RNA to proteins smith2023largescale ; liwang2024protein ; icml2024vqdna ; liu2024genbench . These endeavors aim to leverage deep learning to uncover functional elements, predict gene regulation, and accelerate protein engineering. Notably, the successful application of Transformer-based architectures in natural language processing (NLP) vaswani2017attention ; devlin2019bert has inspired numerous efforts to adapt these methods to biology, giving rise to models for variant effect prediction zhou2022transformer , protein structure elucidation jumper2021alphafold , and more. Despite the promise of these paradigms of foundation models in computational biology lin2023esm ; nguyen2024evo ; shen2024rnafm , three major obstacles continue to hamper multi-omics modeling: (1) Data Island. Current approaches often focus on one molecular modality, modeling only DNA, RNA, or only protein—leading to siloed representations that cannot exploit the inherent relationships among DNA, RNA, and proteins hong2022crossomics , as shown in Figure 1. As a result, critical cross-omics information (for example, how a genetic variant in DNA might alter RNA splicing and ultimately impact protein function) remains underutilized. (2) Naïve “Copy and Paste”. While Transformers vaswani2017attention ; devlin2019bert have revolutionized NLP, simply applying them without incorporating biological knowledge can undermine performance and interpretability rives2021biological . For instance, ignoring codons, coding, and non-coding regions may limit model accuracy and obscure how predictions map to real biological processes. (3) Inefficiency for Long Sequences. Genomic and transcriptomic sequences can span thousands to millions of bases, and mapping proteins back to nucleotide triplets multiplies sequence length alquraishi2019proteinnet . This expansion quickly renders the $\mathcal{O}(n^{2})$ complexity of vanilla Transformers computationally prohibitive, especially for large-scale training on full genomes.
Figure 2: Illustration of Life-Code Framework, which contains three pipelines. (a) Data Pre-processing for the unified input with the central dogma. (b) Pre-training of the Tokenizer and the hybrid Encoder to model the contextual information and the central dogma rules. (c) Transfer Learning to multi-omics downstream tasks with parameter-efficient Supervised Fine-tuning (SFT).
@@ 2.1 Preliminaries
DNA has three notable properties: (1) Length: A chromosome-scale DNA sequence $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$ can reach millions (or more) of base pairs (bp). (2) Double-Helix Complementarity: Each position $x_{i}\in\{\texttt{A},\texttt{T},\texttt{C},\texttt{G}\}$ pairs with a complementary base on the opposite strand. (3) Distributed Regulatory Elements: Important functional sites (e.g., promoters, enhancers, splicing signals) can be located in different regions (CDS or nCDS). In summary, while Transformer-based architectures exhibit strong representational power, their attention mechanism scales with $\mathcal{O}(n^{2})$ complexity, making it difficult to handle sequences of length $n\gg 10^{4}$. Linear or sparse approximations (e.g., Mamba) reduce complexity but often sacrifice representational fidelity for complex biological contexts. Hence, the recent efficient hybrid strategy yang2025deltanet ; yang2024gateddelta ; iclr2024MogaNet ; icml2024chela ; ijcai2024longvq should be considered. We now formally describe how we unify DNA, RNA, and protein sources into nucleotides.
@@ 2.2 Data Pipeline for Life-Code
Figure 3: Data Sampling Pipeline. During pre-training, we sample the DNA sequences and the pair of CDS and amino acids from the DNA dataset and the DNA-AA pairing dataset. Since the DNA sequence sampled from the RefSeq is much longer than the CDS, we predefined the max sequence length (e.g., 8k) for each sample. As for the CDS and its corresponding amino acids, we apply the packing strategy Warner2024ModernBERT to sample with several CDSs.
@@
Following GenBank benson2012genbank curation, we sample from the RefSeq database, truncating long genomic segments and padding shorter ones. For CDS–Amino Acid pairs, we extract annotated CDS records and their corresponding protein sequences, then pack multiple short segments into a single training sample with special separators. See Figure 3 for a schematic illustration. This approach enables us to efficiently process both longer genomic segments and multiple shorter CDS-AA pairs in a unified framework. By balancing truncation/padding for DNA sequences and packing for CDS segments, we ensure that each batch contains a diverse mixture of long-range genomic context and coding information, thereby maximizing the model’s exposure to various biological signals.
@@ 2.3 Life-Code Tokenizer Training
Consider a raw DNA segment of length $3L$. We first apply a 1-d convolution (kernel size $=3$) together with an “unfold” operation (stride $=3$) to transform each consecutive triplet $(x_{3j+1},x_{3j+2},x_{3j+3})$ into a codon-like representation:
$\mathbf{e}_{j}\;=\;\mathrm{Conv1d}(\,x_{3j+1:3j+3}\,).$
@@
$\mathcal{L}_{\mathrm{MLM}}\;=\;\sum_{j=1}^{L}-\log p\bigl{(}\mathbf{x}_{3j+1:3j+3}\,\big{|}\,\widetilde{\mathbf{x}}\bigr{)},$ (2)
@@ 2.4 Bi-Directional Hybrid Model for Long Sequence
To use the double-helix property as Caduceus (icml2024caduceus, ), we split the embeddings $\mathbf{E}=(\mathbf{e}_{1},\mathbf{e}_{2},\ldots,\mathbf{e}_{n})$ into two parts in the feature dimension:
@@
Inspired by recent advances in EVO-style architectures nguyen2024evo and Gated Delta Networks (GDN) yang2024gateddelta , we design the encoder in Life-Code as the linear-attention Gated DeltaNet mixed with a small proportion of standard multi-head self-attention (MHSA). Concretely, for every twelve encoder layers, eleven employ the GDN update rule—offering linear or near-linear complexity via kernel-based reordering—while one layer applies MHSA. The GDN layer relies on a delta update to maintain memory, reducing the quadratic overhead of naive Transformers. Meanwhile, the one full-attention layer provides additional global context, which is crucial for capturing fine-grained long-range dependencies. This 11 : 1 ratio of GDN to dense layers significantly boosts computational efficiency (compared to a pure $\mathcal{O}(n^{2})$ Transformer) while preserving model expressiveness and long-distance modeling capacity.
@@
For the distillation step, we replace the previous squared-error formulation with a token-wise KL loss on the predicted distributions, which we have empirically found to be more robust (e.g., no negative log issues). Inspired by the DINO framework (oquab2024dinov2, ), we decode the encoder’s final protein sequence embeddings to protein structure embeddings a Protein Decoder $\mathcal{D}$ to produce a per-token distribution aligned with the teacher PLM like ESM2 (lin2022ESM2, ). Formally, let $P_{\theta}(\mathcal{D}(\mathbf{h}_{j}^{\mathrm{LifeCode}}))\in\mathbb{R}^{D^{\prime}}$ and $P_{\theta^{\prime}}(\mathbf{h}_{j}^{\mathrm{PLM}})\in\mathbb{R}^{D^{\prime}}$ be the predicted distributions for token $j$ from our model and the teacher, where the Decoder can be a two-layer-MLP with $D\rightarrow 4D\rightarrow D^{\prime}$. Then, the alignment or KD loss is defined as:
@@
We combine these three losses using a log-sum multi-task strategy (lin2023dualmtl, ) so that all terms (MLM, Translation, and KD) are treated as entropy-like objectives:
$\log\bigl{(}\mathcal{L}_{\mathrm{MLM}}+\epsilon\bigr{)}\;+\;\log\bigl{(}\mathcal{L}_{\mathrm{Trans}}+\epsilon\bigr{)}\;+\;\log\bigl{(}\mathcal{L}_{\mathrm{KD}}+\epsilon\bigr{)},$ (5)
@@
Figure 5: Illustration of Life-Code Encoder pre-training pipeline. We model the long sequence by the hybrid encoder as a mixture of RNN blocks (DeltaNet) and self-attention blocks. With pre-trained tokenizers and de-tokenizers, we consider three types of pre-training tasks: (a) the nucleotide masked modeling for contextual information, (b) the cDNA-AA translation for the central dogma (with span masks), and (c) the knowledge distillation (KD) with pre-trained protein LM. Note that red lines indicates the learning objectives.
Figure 6: Empirical Analysis of Life-Code Tokenizer. Left: Codon Usage Bias in Life-Code Tokenizer across four representative species—E. coli, S. cerevisiae (yeast), D. melanogaster (fruit fly), and H. sapiens (human)—illustrating variations in codon frequency (%) for amino acids. The pastel strips highlight codons belonging to the same amino acid group. Right: t-SNE visualization of learned codon embeddings in Life-Code Tokenizer, where codons that translate to the same or biochemically similar amino acids cluster together (e.g., hydrophobic or charged groups), and stop codons (black crosses) form a distinct region. These views jointly demonstrate that Life-Code encodes both species-dependent codon usage preferences and biologically meaningful relationships.
@@ 2.5 Empirical Consequences
In Figure 6, our analyses reveal that the Life-Code Tokenizer captures biologically meaningful codon relationships in two complementary ways. First, codon usage bias (left figure) indicates that different species, from E. coli to H. sapiens, vary markedly in their preference for certain codons—an effect traditionally tied to tRNA abundance and translational efficiency. By visualizing these preferences across amino acids, we observe how each organism clusters around specific high-frequency codons, underscoring the ecological and evolutionary factors shaping codon selection. Second, the t-SNE plot of the tokenizer’s learned codon embeddings (right figure) demonstrates that codons translating to amino acids with similar biochemical properties (e.g., hydrophobic or charged) are embedded close together, while the stop codons occupy a separate region. This structural arrangement confirms that the model internalizes codon–amino acid mappings, reflecting both frequency patterns and higher-level attributes like polarity or aromaticity. Together, these findings suggest that the Life-Code Tokenizer not only memorizes sequence data but also discerns and encodes intrinsic biological signals—ranging from species-specific usage preferences.
@@
We evaluated the throughput (thousands of tokens processed per second) of Life-Code, EVO, and DNABERT-2 on a single A100 GPU under increasing sequence lengths (with batch size adjusted accordingly). As shown in Figure 7, DNABERT-2 (blue) exhibits a steep drop-off at longer sequences and ultimately runs out of memory (OOM) at $16\text{K}\times 2$. In contrast, EVO (red) and Life-Code (green) both maintain higher throughput thanks to their hybrid attention mechanisms. Notably, Life-Code achieves the best scalability, retaining above 40K t/s even at $8\text{K}\times 4$. These results confirm that Life-Code’s efficient architecture can accommodate extensive input lengths with minimal degradation in training speed, making it well-suited for large-scale genomic modeling.
@@ 3.1 Experimental Setup
With the nucleotide vocabulary, the Life-Code tokenizer contains the following modules: (i) A linear projection from the DNA input to 384-dim implemented by nn.Embedding; (ii) A GDN block (DeltaNet) with 384-dim for global contextual modeling; (iii) A 1-d Convolution with a kernel size of 3 and stride of 1, followed by an UnFold operation to merge every three nucleotide tokens into 768-dim codon embedding. Similarly, we also design the DNA De-Tokenizer and Amino Acid Translator for tokenizer pre-training. As shown in Figure 4, we first pre-train the Life-Code tokenizer by AdamW optimizer (iclr2019AdamW, ) for 100,000 iterations with a basic learning rate of $2\times 10^{-4}$ adjusted by the Cosine scheduler and a total batch size of 512, as summarized in Table A2. As shown in Table A1, the two pre-training datasets (DNA and DNA-AA pairing) are constructed from NCBI, GenBank, and UniRef50 databases, as described in Figure A1, Appendix A, and  A.2.
@@
Similar to the BERT-Large architecture (devlin2019bert, ), the Life-Code encoder has 24 layers in total with an embedding dim of 1024/1280 and 340/500M parameters. We enhance the model with three-fold designs: (i) Mixture of DeltaNet (yang2025deltanet, ) blocks and Multi-head Self-attention (MHSA) blocks, especially every 11 DeltaNet blocks followed by one MHSA block. (ii) The LLaMA-like macro design (Touvron2023LLaMA, ) with RMSNorm (Zhang2019RMSNorm, ), Layer Scale (iccv2021CaiT, ), Rotary Position Embedding (RoPE) (Su2021RoFormer, ), SwiGLU, and FlashAttention to facilitate stable pre-training with long sequences. (iii) During pre-training, we apply the packing strategy (Warner2024ModernBERT, ) to build up a long sequence with several CDS, which compromises the gap between different lengths of the reference sequence and the coding sequences. With three pre-training tasks in Eq. 5, the Life-Code tokenizer and encoder are optimized by the AdamW optimizer for 1M iterations with a batch size of 256 and a basic learning rate of $1\times 10^{-4}$. We adopt 15% random masking in BERT for Masked DNA Reconstruction and the 3-mer span masking for CDS-to-Amino-Acid Translation. As for KD from a pre-trained Protein LM, we adopted ESM2-650M (lin2022ESM2, ) and a protein decoder with an output dimension of 1280. During the warmup periods, the maximum sequence length is 1024, with a linear warmup of the learning rate for 10 iterations. After that, the maximum sequence length is set to 4k with the learning rate adjusted by the Cosine Annealing scheduler. View Appendix A.2 for more details.
@@
Average 87.07 85.02 87.30 87.94 89.26 85.89 87.69 89.12 86.95 90.79
Table 2: Comparison on GUE Benchmark. MCC (%) averaged across several sub-tasks is reported with SFT evaluation for seven primary tasks (24 sub-tasks in total). The best and the second best results are marked as the bold and underlined types.
@@
Human TF Detection (3) -- 64.17 63.34 70.11 -- 70.56 -- -- 71.34
Mouse TF Detection (3) -- 56.43 67.02 67.99 -- 69.80 -- -- 71.76
Core Promoter Detection (3) -- 71.81 71.63 70.53 -- 73.37 -- -- 74.53
Promoter Detection (3) -- 81.69 88.15 84.21 -- 86.58 -- -- 89.28
Splice Site Reconstructed (1) -- 84.07 89.35 84.99 -- 89.53 -- -- 90.24
Virus Covid Classification (1) -- 55.50 73.04 71.02 -- 74.32 -- -- 74.88

### v3

@@ 1 Introduction
The advent of large-scale biological datasets has sparked extensive interest in sequence modeling across multiple omics domains—ranging from DNA and RNA to proteins (liwang2024protein; icml2024vqdna; liu2024genbench). These endeavors aim to leverage deep learning to uncover functional elements, predict gene regulation, and accelerate protein engineering. Notably, the successful application of Transformer-based architectures in natural language processing (NLP) (vaswani2017attention; devlin2019bert) has inspired numerous efforts to adapt these methods to biology, giving rise to models for variant effect prediction (zhou2022transformer), protein structure elucidation (jumper2021alphafold), and more. Despite the promise of these paradigms of foundation models in computational biology (lin2023esm; nguyen2024evo; shen2024rnafm), three major obstacles continue to hamper multi-omics modeling: (1) Data Island. Current approaches often focus on one molecular modality, modeling only DNA, RNA, or only protein—leading to siloed representations that cannot exploit the inherent relationships among DNA, RNA, and proteins (hong2022crossomics), as shown in Figure 1. As a result, critical cross-omics information (for example, how a genetic variant in DNA might alter RNA splicing and ultimately impact protein function) remains underutilized. (2) Naïve “Copy and Paste”. While Transformers (vaswani2017attention; devlin2019bert) have revolutionized NLP, simply applying them without incorporating biological knowledge can undermine performance and interpretability (rives2021biological). For instance, ignoring codons, coding, and non-coding regions may limit model accuracy and obscure how predictions map to real biological processes. (3) Inefficiency for Long Sequences. Genomic and transcriptomic sequences can span thousands to millions of bases, and mapping proteins back to nucleotide triplets multiplies sequence length (alquraishi2019proteinnet). This expansion quickly renders the $\mathcal{O}(n^{2})$ complexity of vanilla Transformers computationally prohibitive, especially for large-scale training on full genomes.
Figure 1: Overview of the Central Dogma Data Flow with biological language models (LM), separately proposed.
Figure 2: Illustration of Life-Code Framework, which contains three pipelines. (a) Data Pre-processing for the unified input with the central dogma. (b) Pre-training of the Tokenizer and the hybrid Encoder to model the contextual information and the central dogma rules. (c) Transfer Learning to multi-omics downstream tasks with parameter-efficient Supervised Fine-tuning.
@@ 2.1 Preliminaries
DNA has three notable properties: (1) Length: A chromosome-scale DNA sequence $\mathbf{x}=(x_{1},x_{2},\ldots,x_{n})$ can reach millions (or more) of base pairs (bp). (2) Double-Helix Complementarity: Each position $x_{i}\in\{\texttt{A},\texttt{T},\texttt{C},\texttt{G}\}$ pairs with a complementary base on the opposite strand. (3) Distributed Regulatory Elements: Important functional sites (e.g., promoters, enhancers, splicing signals) can be located in different regions (CDS or nCDS). In summary, while Transformer-based architectures exhibit strong representational power, their attention mechanism scales with $\mathcal{O}(n^{2})$ complexity, making it difficult to handle sequences of length $n\gg 10^{4}$. Linear or sparse approximations (e.g., Mamba) reduce complexity but often sacrifice representational fidelity for complex biological contexts. Hence, the recent efficient hybrid strategy (yang2025deltanet; yang2024gateddelta; iclr2024MogaNet; icml2024chela; ijcai2024longvq) should be considered. We now formally describe how we unify DNA, RNA, and protein sources into nucleotides.
@@ 2.2 Data Pipeline for Life-Code
Figure 3: Data Sampling Pipeline. During pre-training, we sample the DNA sequences and the pair of CDS and amino acids from the DNA dataset and the DNA-AA pairing dataset. Since the DNA sequence sampled from the RefSeq is much longer than the CDS, we predefined the max sequence length (e.g., 8k) for each sample. As for the CDS and its corresponding amino acids, we apply the packing strategy (Warner2024ModernBERT) to sample with several CDSs.
@@
Following GenBank (benson2012genbank) curation, we sample from the RefSeq database, truncating long genomic segments and padding shorter ones. For CDS–Amino Acid pairs, we extract annotated CDS records and their corresponding protein sequences, then pack multiple short segments into a single training sample with special separators. See Figure 3 for a schematic illustration. This approach enables us to efficiently process both longer genomic segments and multiple shorter CDS-AA pairs in a unified framework. By balancing truncation/padding for DNA sequences and packing for CDS segments, we ensure that each batch contains a diverse mixture of long-range genomic context and coding information, thereby maximizing the model’s exposure to various biological signals.
@@ 2.3 Life-Code Tokenizer Training
Consider a raw DNA segment of length $3L$. We first apply a 1-d convolution (kernel size =3=3) together with an “unfold” operation (stride =3=3) to transform each consecutive triplet $(x_{3j+1},x_{3j+2},x_{3j+3})$ into a codon-like representation:
$\mathbf{e}_{j}\;=\;\mathrm{Conv1d}(\,\mathbf{x}_{3j+1:3j+3}\,).$
@@
[…]
@@ 2.4 Bi-Directional Hybrid Model for Long Sequence
To use the double-helix property as Caduceus (icml2024caduceus), we split the embeddings $\mathbf{E}=(\mathbf{e}_{1},\mathbf{e}_{2},\ldots,\mathbf{e}_{n})$ into two parts in the feature dimension:
@@
Figure 5: Empirical Analysis of Life-Code Tokenizer. Left: Codon Usage Bias in Life-Code Tokenizer across four representative species—E. coli, S. cerevisiae (yeast), D. melanogaster (fruit fly), and H. sapiens (human)—illustrating variations in codon frequency (%) for amino acids. The pastel strips highlight codons belonging to the same amino acid group. Right: t-SNE visualization of learned codon embeddings in Life-Code Tokenizer, where codons that translate to the same or biochemically similar amino acids cluster together (e.g., hydrophobic or charged groups), and stop codons form a distinct region.
@@
Inspired by recent advances in EVO-style architectures (nguyen2024evo) and Gated Delta Networks (GDN) (yang2024gateddelta), we design the encoder in Life-Code as the linear-attention Gated DeltaNet mixed with a small proportion of standard multi-head self-attention (MHSA). Concretely, for every twelve encoder layers, eleven employ the GDN update rule—offering linear or near-linear complexity via kernel-based reordering—while one layer applies MHSA. The GDN layer relies on a delta update to maintain memory, reducing the quadratic overhead of naive Transformers. Meanwhile, the one full-attention layer provides additional global context, which is crucial for capturing fine-grained long-range dependencies. This 11 : 1 ratio of GDN to dense layers significantly boosts computational efficiency (compared to a pure $\mathcal{O}(n^{2})$ Transformer) while preserving model expressiveness and long-distance modeling capacity.
@@
For the distillation step, we replace the previous squared-error formulation with a token-wise KL loss on the predicted distributions, which we have empirically found to be more robust (e.g., no negative log issues). Inspired by the DINO framework (oquab2024dinov2), we decode the encoder’s final protein sequence embeddings to protein structure embeddings a Protein Decoder $\mathcal{D}$ to produce a per-token distribution aligned with the teacher PLM like ESM2 (lin2022ESM2). Formally, let $P_{\theta}(\mathcal{D}(\mathbf{h}_{j}^{\mathrm{LifeCode}}))\in\mathbb{R}^{D^{\prime}}$ and $P_{\theta^{\prime}}(\mathbf{h}_{j}^{\mathrm{PLM}})\in\mathbb{R}^{D^{\prime}}$ be the predicted distributions for token $j$ from our model and the teacher, where the Decoder can be a two-layer-MLP with $D\rightarrow 4D\rightarrow D^{\prime}$. Then, the alignment or KD loss is defined as:
@@
We combine these three losses using a log-sum multi-task strategy (lin2023dualmtl) so that all terms (MLM, Translation, and KD) are treated as entropy-like objectives:
$\hskip-10.00002pt\log\bigl(\mathcal{L}_{\mathrm{MLM}}+\epsilon\bigr)\;+\;\log\bigl(\mathcal{L}_{\mathrm{Trans}}+\epsilon\bigr)\;+\;\log\bigl(\mathcal{L}_{\mathrm{KD}}+\epsilon\bigr),$ (5)
@@
Figure 7: Illustration of Life-Code Encoder pre-training pipeline. We model the long sequence by the hybrid encoder as a mixture of DeltaNet and Attention blocks. With pre-trained tokenizers and de-tokenizers, we consider three types of pre-training tasks: (a) the nucleotide masked modeling for contextual information, (b) the cDNA-AA translation for the central dogma, and (c) the knowledge distillation (KD) with pre-trained protein LM. Note that red lines indicates the learning objectives.
@@ 2.5 Empirical Consequences
In Figure 5, our analyses reveal that the Life-Code Tokenizer captures biologically meaningful codon relationships in two complementary ways. First, codon usage bias (left figure) indicates that different species, from E. coli to H. sapiens, vary markedly in their preference for certain codons—an effect traditionally tied to tRNA abundance and translational efficiency. By visualizing these preferences across amino acids, we observe how each organism clusters around specific high-frequency codons, underscoring the ecological and evolutionary factors shaping codon selection. Second, the t-SNE plot of the tokenizer’s learned codon embeddings (right figure) demonstrates that codons translating to amino acids with similar biochemical properties (e.g., hydrophobic or charged) are embedded close together, while the stop codons occupy a separate region. This structural arrangement confirms that the model internalizes codon–amino acid mappings, reflecting both frequency patterns and higher-level attributes like polarity or aromaticity. Together, these findings suggest that the Life-Code Tokenizer not only memorizes sequence data but also discerns and encodes intrinsic biological signals.
@@
We evaluated the throughput (thousands of tokens processed per second) of Life-Code, EVO, and DNABERT-2 on a single A100 GPU under increasing sequence lengths (with batch size adjusted accordingly). As shown in Figure 6, DNABERT-2 (blue) exhibits a steep drop-off at longer sequences and ultimately runs out of memory (OOM) at $16\text{K}\times 2$. In contrast, EVO (red) and Life-Code (green) both maintain higher throughput thanks to their hybrid attention mechanisms. Notably, Life-Code achieves the best scalability, retaining above 40K t/s even at $8\text{K}\times 4$. These results confirm that Life-Code’s efficient architecture can accommodate extensive input lengths with minimal degradation in training speed.
@@ 3.1 Experimental Setup
With the nucleotide vocabulary, the Life-Code tokenizer contains the following modules: (i) A linear projection from the DNA input to 384-dim implemented by nn.Embedding; (ii) A GDN block (DeltaNet) with 384-dim for global contextual modeling; (iii) A 1-d Convolution with a kernel size of 3 and stride of 1, followed by an UnFold operation to merge every three nucleotide tokens into 768-dim codon embedding. Similarly, we also design the DNA De-Tokenizer and Amino Acid Translator for tokenizer pre-training. As shown in Figure 4, we first pre-train the Life-Code tokenizer by AdamW optimizer (iclr2019AdamW) for 100,000 iterations with a basic learning rate of $2\times 10^{-4}$ adjusted by the Cosine scheduler and a total batch size of 512, as summarized in Table A2. As shown in Table A1 and  A1, the two pre-training datasets (DNA and DNA-AA pairing) are constructed from NCBI, GenBank, and UniRef50 databases.
@@
Similar to the BERT-Large architecture (devlin2019bert), the Life-Code encoder has 24 layers in total with an embedding dim of 1024/1280 and 340/500M parameters. We enhance the model with three-fold designs: (i) Mixture of DeltaNet (yang2025deltanet) blocks and Multi-head Self-attention (MHSA) blocks, especially every 11 DeltaNet blocks followed by one MHSA block. (ii) The LLaMA-like macro design (Touvron2023LLaMA) with RMSNorm (Zhang2019RMSNorm), Layer Scale (iccv2021CaiT), Rotary Position Embedding (RoPE) (Su2021RoFormer), SwiGLU, and FlashAttention to facilitate stable pre-training with long sequences. (iii) During pre-training, we apply the packing strategy (Warner2024ModernBERT) to build up a long sequence with several CDS, which compromises the gap between different lengths of the reference sequence and the coding sequences. With three pre-training tasks in Eq. 5, the Life-Code tokenizer and encoder are optimized by the AdamW optimizer for 1M iterations with a batch size of 256 and a basic learning rate of $1\times 10^{-4}$. We adopt 15% random masking in BERT for Masked DNA Reconstruction and the 3-mer span masking for CDS-to-Amino-Acid Translation. As for KD from a pre-trained Protein LM, we adopted ESM2-650M (lin2022ESM2) and a protein decoder with an output dimension of 1280. During the warmup periods, the maximum sequence length is 1024, with a linear warmup of the learning rate for 10 iterations. After that, the maximum sequence length is set to 4k with the learning rate adjusted by the Cosine Annealing scheduler. View Appendix A.2 for more details.
@@
Average (8 tasks) 87.07 85.02 87.30 87.94 89.26 85.89 87.69 89.12 86.95 90.79
@@
Human TF Detection (3) - 64.17 63.34 70.11 - 70.56 - - 71.34
Mouse TF Detection (3) - 56.43 67.02 67.99 - 69.80 - - 71.76
Core Promoter Detection (3) - 71.81 71.63 70.53 - 73.37 - - 74.53
Promoter Detection (3) - 81.69 88.15 84.21 - 86.58 - - 89.28
Splice Site Reconstructed (1) - 84.07 89.35 84.99 - 89.53 - - 90.24
Virus Covid Classification (1) - 55.50 73.04 71.02 - 74.32 - - 74.88
@@
Table 2: Comparison on GUE Benchmark. Matthews Correlation Coefficient (MCC) (%) averaged across several sub-tasks is reported with SFT evaluation. The best and the second best results are marked as the bold and underlined types.